# Bloom: Designing for LLM-Augmented Behavior Change Interactions

### v1

@@ 5.3. Procedure
Physical Activity & Health (Custom) General health self-assessment, satisfaction with PA & health, and motivation to change (custom questions). All questions are provided in Appendix Table 8.
@@
Barriers to Being Active Quiz (Rodríguez et al., 2008) [^6] https://www.cdc.gov/diabetes/professional-info/pdfs/toolkits/road-to-health-barriers-activity-quiz-p.pdf Likelihood that common barriers to regular PA apply to oneself.
User Experience & Advice Quality (adapted from (Jörke et al., 2025)) Subjective experience with the system and quality of advice received. All questions are provided in Appendix Table 8.
@@
System Usability Scale (SUS) (Brooke et al., 1996) General system usability.
@@
5.4. Measures
5.4.1. Wearable Data
Participants consented to share PA data from HealthKit, including step count, distance walking/running, active energy burned, exercise time, workouts, and heart rate. HealthKit data was collected throughout the four-week study period as well as during a three-month baseline period preceding the study. Participants were informed during onboarding that they could disable any data source they were uncomfortable sharing at any time via built-in system privacy settings.
@@ 5.4.2. Survey Data
Participants completed onboarding (pre-study), offboarding (post-study), weekly, and (optional) daily surveys. Onboarding and offboarding surveys are summarized in Table 6. Weekly surveys included Likert-scale and free-response questions regarding participants’ PA, health, user experience, a four-item subset of Process Mindset, and a subset of Barriers to Being Active. Daily surveys consisted of five brief Likert-scale questions regarding participants’ PA, health hopefulness, mood, and an optional free-response reflection.
@@ 5.5.1. Qualitative Coding
We performed qualitative coding on participants’ onboarding and offboarding interview transcripts using thematic analysis (Braun and Clarke, 2012). Two authors collaboratively coded two participants and one author coded five participants, after which an initial codeset was created. One author coded the remaining transcripts, with frequent meetings to discuss codes, emerging themes, and differences in interpretation.
@@ 3. H3:
We analyzed each outcome measure using a three-level linear mixed-effects models in which daily observations (level 1) are nested within study weeks (level 2), nested within participants (level 3). Fixed effects include study period (baseline vs. study), treatment condition (LLM vs. control), study week (1–4), and day-of-week. Random intercepts at the participant and participant-week level account for repeated measurements and within-subject variability. This modeling approach was selected as it provided the best fit to our data and required less stringent assumptions about linearity of treatment effects. Further details are provided in Appendix C.1.
@@ 6. Results
In this section, we report on our study’s findings using mixed methods. In our semi-structured interviews, participants described Beebo’s conversational style as supportive and non-prescriptive, which many linked to more positive and nuanced shifts in their own mindsets towards PA. At the same time, they most often cited plans, notifications, and the ambient display—not the chat itself—as their primary sources of accountability.
We observed significant increases in PA levels measured with wearable data relative to baseline across both conditions, but did not find statistically significant treatment–control differences. App usage and plan data indicated greater engagement and personalization in the treatment condition: participants spent over five times longer in the app than controls, with increased usage across all screens (not just chat), and produced more varied plans with slightly higher completion rates.
@@ 6.1. Qualitative Coding
Participants in both groups reported positive experiences with the application and benefits to their PA habits, but the groups differed substantially in the language used to describe the application and changes in their attitudes towards PA. Below, we summarize the key themes derived from qualitative coding of participant interviews.
@@ 6.1.2. Beebo Supported Accountability Through Flexible Planning and Personalized Advice
Notably, participants rarely identified Beebo itself as the single most useful feature, but rather described it as valuable in supporting and reinforcing the accountability provided by other app features. For example, participants appreciated that Beebo could incorporate personal preferences into their weekly plans, such as P23: “It was nice that Beebo […] would take into account me telling him, plan it Tuesday, Thursdays in the morning, Wednesday, Fridays in the afternoon.” Others described how collaborative goal-setting reinforced motivation, such as P22: “I would say it’s a real motivation, having the goal set up versus just using, like, the Apple Fitness goals. Having that interactivity between the AI and the chat was another motivational goal that helped help me be successful.”
@@
Taken together, these results highlight that while the chatbot itself was not typically seen as the central source of accountability or motivation, participants valued the supportive and facilitative role it played in creating and flexibly adjusting weekly plans, overcoming unforeseen barriers, and providing tailored recommendations and advice.
@@ 6.1.3. Positive and Supportive Tone Mirrored Positive Shifts in Physical Activity Mindsets
Treatment participants consistently described Beebo’s conversational style in empathetic, supportive, non-judg-mental, and collaborative terms. P34 explained, “When I used to go to the gym, I felt like I was pushed all the time by my trainer. […] Beebo doesn’t do that. […] Beebo is more gentle and more my pace for my age, anyway. I always felt overwhelmed. With Beebo, I feel like I’m supported” (P34). Participants explicitly highlighted Beebo’s gentle approach as critical to their positive shifts in mindset: “I thought it was going to be more difficult. […] Instead, it was working at my own pace […] asking, ’How’s it feeling for you?’ […] I liked the gentler approach” (P6). Despite full awareness that Beebo is an AI chatbot, participants often explicitly described Beebo as being empathetic, such as P52, “I wouldn’t say personal because it’s not real, but like more of an empathetic, encouraging type of tone of voice.”
@@
While participants in both conditions reported changes in their attitudes towards PA, treatment participants articulated changes in their perceived benefits of PA with greater specificity and nuance. Control participants spoke about changes in their mindset mostly as a result of having gotten more exercise. “I don’t think it changed the way I feel about physical activity other than being more motivated to do it” (P15).
Treatment participants mentioned a far greater diversity of changes, including greater self-confidence, more self-forgiveness around missed activities, expanded views on what constitutes PA, and increased intrinsic enjoyment of exercise. For instance, P13 expressed a greater appreciation for everyday physical activities: “My attitude going into this was, well, I don’t do a lot, I’m not doing enough. […] And this helped me understand that I actually am doing things. When I work out in the garden and I’m digging holes, […] I didn’t think that it was exercise or fitness in any way.” This increase in awareness led P13 to experience a gradual shift in confidence: “Each day I got a little bit more confident. […] It wasn’t a light bulb moment. It was a small, everyday, slow build.” Similarly, P34 linked increased awareness of movement directly to their interactions with Beebo: “It definitely has changed the way I think about it […] if I go into the kitchen or outside or something, I’m doing all these additional steps. […] I’m moving a whole lot more than I ever did. And I think it’s because of my little Beebo.” Lastly, some participants described an increased intrinsic enjoyment of PA, such as P1: “At first I was like, working out sucks. But now I’m kind of like, oh, no. It’s kind of just part of what I do and I enjoy. I look forward to working out now.”
@@ 6.1.4. Opportunities to Increase Accountability, Specificity, and Diversity
Figure 7. Mindset-related survey outcomes. We plot mindset-related survey items for pre/post (top), weekly (middle), and daily (bottom) survey responses. Scores represent group means on a 5-point Likert scale, with higher values reflecting more positive orientations toward PA. Across various mindset-related survey measures, treatment participants showed greater increases than control participants.
@@ 6.2. Survey Data
Next, we report on survey outcomes organized by topic: physical activity, mindset and satisfaction, usability, and user experience. As discussed in Section 5.5.2, no statistical significance testing was conducted and findings should be interpreted as descriptive patterns of change. Full survey results are provided in Appendix Tables 9 (pre/post), 10 (daily), and 11 (weekly).
@@ 6.2.1. Physical Activity
Self-reported levels of PA, assessed via the International Physical Activity Questionnaire (IPAQ), increased in both groups with similar magnitude. Exercise Stage of Change (1 = pre-contemplation, 5 = action) advanced on average by +1.0+1.0 stages in the treatment group and by +0.7+0.7 in the control group.
@@ 6.2.2. Mindset & Satisfaction
Mindset-related outcomes exhibited the most prominent treatment-control differences, as shown in Figure 7. Treatment participants reported a larger increase in PA adequacy mindset, namely the belief that PA is beneficial to their health (+1.2+1.2 points from pre- to post-study vs. +0.8+0.8 in control; 5-point Likert scale), while both groups showed a similar increases in process mindset, the degree to which exercise is seen as inherently enjoyable (+0.3+0.3 vs. +0.2+0.2 points). Satisfaction with PA levels showed larger increases in the treatment group (+1.4+1.4 vs. +1.1+1.1 points), as did satisfaction with physical health (+0.9+0.9 vs. +0.6+0.6 points). Meanwhile, perceived barriers to activity declined slightly in both groups (0.2-0.2 vs. 0.2-0.2 points), while self-efficacy was stable ($0.0$ vs. $0.0$ points), suggesting that the groups differed in their beliefs about the benefits and enjoyment of PA, but not in their perceived barriers or efficacy.
Weekly and daily surveys exhibit similar trends. Treatment participants showed more positive slopes in satisfaction with their current PA levels (+0.02+0.02 vs. +0.01+0.01 points/day), identifying as a physically active person (+0.03+0.03 vs. +0.01+0.01 points/day), and hopefulness about their health (+0.01+0.01 vs. $0.00$ points/day). Treatment participants also showed greater increases in perceptions that exercising is easy (+0.1+0.1 vs. +0.0+0.0 points/week; on a scale from 1: difficult to 5: easy).
@@
Burned (kcal) H2 Treatment–control difference in daily kcal burned change from baseline to study period –7.09 (31.3)
@@
Running (km) H2 Treatment–control difference in daily distance walked/run (km) change from baseline to study period –0.239 (0.198)
H3 Treatment–control difference in the weekly rate of change ($\Delta$ km/day per study week) +0.071 (0.063)
Table 7. Wearable Data (Quantitative) Results. We report mean (SE) parameter estimates from our model for each hypothesis. *** denotes $p<0.001$ and ** denotes $p<0.01$ ($p$-values are Holm‑adjusted for multiple comparisons). Significant results are bolded.
6.2.3. Usability
@@ 6.2.4. User Experience
Participants in the treatment group consistently rated the quality of advice and overall user experience more favorably than the control group at both pre- and post-study measurements. We replicate Jörke et al.’s (Jörke et al., 2025) strong initial impressions of the GPTCoach system (pre-study mean of $4.4$ in treatment vs. $4.1$ in control; 5-point Likert scale). However, as with usability, ratings declined pre- to post-study in both conditions (0.4-0.4 vs. 0.5-0.5 points).
Treatment participants rated shared decision-making more highly at both pre- and post-study measurements, with scores increasing pre/post in treatment while decreasing in control (+0.1+0.1 vs. 0.2-0.2; 5-point Likert scale), suggesting that goal setting was perceived as a more collaborative activity in the treatment group. The treatment group also reported greater increases in insights into their personal data, measured via the Technology Support Reflection Inventory (TSRI) than the control group (+0.1+0.1 vs. 0.3-0.3; 5-point Likert scale). In addition, treatment participants also rated their interactions as more human-like (post-study mean of $3.5$ vs. $2.83$).

### v2

@@ 5.4.1. Wearable Data
Physical Activity & Health (Custom) General health self-assessment, satisfaction with PA & health, and motivation to change (custom questions). All questions are provided in Supplementary Table 10.
@@
Barriers to Being Active Quiz (Centers for Disease Control and Prevention, 2022) [^6] https://www.cdc.gov/diabetes/professional-info/pdfs/toolkits/road-to-health-barriers-activity-quiz-p.pdf Likelihood that common barriers to regular PA apply to oneself.
User Experience & Advice Quality (adapted from (Jörke et al., 2025)) Subjective experience with the system and quality of advice received. All questions are provided in Supplementary Table 10.
@@
System Usability Scale (SUS) (Brooke and others, 1996) General system usability.
@@ 5.4.2. Survey Data
Participants completed onboarding (pre-study), offboarding (post-study), weekly, and (optional) daily surveys. Onboarding and offboarding surveys are summarized in Table 6. Weekly surveys included Likert scale and free response questions regarding participants’ PA, health, user experience, a four-item subset of Process Mindset, and a subset of Barriers to Being Active. Daily surveys consisted of five brief Likert scale questions regarding participants’ PA, health hopefulness, mood, and an optional free response reflection. This set of survey measures was sourced from prior work on PA promotion in HCI (Murnane et al., 2020; Jörke et al., 2025; Sefidgar et al., 2024; Mitchell et al., 2021), and aimed to have broad coverage of relevant PA constructs.
@@ 5.5.1. Qualitative Coding
We performed qualitative coding on participants’ onboarding and offboarding interview transcripts using reflexive thematic analysis (Braun and Clarke, 2006, 2019). Two authors first collaboratively reviewed and coded several transcripts to identify important aspects of data to attend to. One author coded the remaining transcripts, generating and iteratively developing codes as new insights arose. Throughout the analysis, both researchers met regularly to reflect on and discuss similarities and differences in their interpretations of the data.
@@ 3. H3:
H1 measures whether the study produced meaningful improvements over the pre-study baseline across both conditions. This serves as a validation check that the study period elicited meaningful behavior change before interpreting between-condition differences. H2 and H3 test for differences between the treatment and control groups for mean PA levels (H2) and PA persistence (H3) during the study period. While our sample size provides adequate power for H1, the between-condition comparisons in H2 and H3 are likely underpowered for detecting small to moderate effects (see Supplementary Material C.2). We therefore report results for H2 and H3 as exploratory analyses to transparently characterize directional patterns that may guide future, fully powered trials.
We analyzed each outcome measure using a three-level linear mixed-effects model in which daily observations (level 1) are nested within study weeks (level 2), nested within participants (level 3). Fixed effects include study period (baseline vs. study), treatment condition (LLM vs. control), study week (1–4), and day-of-week. Random intercepts at the participant and participant-week level account for repeated measurements and within-subject variability. This modeling approach was selected as it provided the best fit to our data and required less stringent assumptions about linearity of treatment effects. Further details are provided in Supplementary Material C.1.
@@ 6. Results
In this section, we report on the findings from our mixed-methods evaluation. We first report on qualitative coding of our semi-structured interviews, followed by survey outcomes. We then present statistical analyses of wearable PA outcomes and conclude with plan and app usage data. Our study’s findings are summarized in Table 8.
@@ 6.1. Qualitative Coding
Participants in both groups reported positive experiences with the application and benefits to their PA habits, but the groups differed substantially in the language used to describe changes in their beliefs about PA. Below, we summarize the key themes derived from qualitative coding of participant interviews.
@@ 6.1.2. Beebo Supported Accountability Through Flexible Planning and Personalized Advice
Notably, participants rarely identified Beebo itself as the single most useful feature, but rather described it as valuable in supporting and reinforcing the accountability provided by other app features. For example, participants appreciated that Beebo could incorporate personal preferences into their weekly plans, such as P23: “It was nice that Beebo […] would take into account me telling him, plan it Tuesday, Thursdays in the morning, Wednesday, Fridays in the afternoon.” Others described how collaborative goal setting reinforced motivation, such as P22: “I would say it’s a real motivation, having the goal set up versus just using, like, the Apple Fitness goals. Having that interactivity between the AI and the chat was another motivational goal that helped help me be successful.”
@@
Taken together, these results highlight that while the chatbot was not seen as the central source of accountability or motivation, participants valued the supportive and facilitative role it played in helping create and flexibly adjust weekly plans, overcome unforeseen barriers, and provide tailored recommendations and advice.
@@ 6.1.3. Positive and Supportive Tone Mirrored Positive Shifts in Physical Activity Mindsets
Treatment participants consistently described Beebo’s conversational style in empathetic, supportive, non-judgmental, and collaborative terms. P34 explained, “When I used to go to the gym, I felt like I was pushed all the time by my trainer. […] Beebo doesn’t do that. […] Beebo is more gentle and more my pace for my age, anyway. I always felt overwhelmed. With Beebo, I feel like I’m supported” (P34). Participants explicitly highlighted Beebo’s gentle approach as critical to their positive shifts in mindset: “I thought it was going to be more difficult. […] Instead, it was working at my own pace […] asking, ’How’s it feeling for you?’ […] I liked the gentler approach” (P6). Despite full awareness that Beebo is an AI chatbot, participants often explicitly described Beebo as being empathetic, such as P52, “I wouldn’t say personal because it’s not real, but like more of an empathetic, encouraging type of tone of voice.”
@@
While participants in both conditions reported changes in their attitudes toward PA, treatment participants articulated changes in their perceived benefits of PA with greater specificity and nuance. Control participants spoke about changes in their mindset mostly as a result of having gotten more exercise. “I don’t think it changed the way I feel about physical activity other than being more motivated to do it” (P15). Meanwhile, treatment participants mentioned a far greater diversity of changes, including greater self-confidence, more self-compassion around missed activities, expanded views on what constitutes PA, and increased intrinsic enjoyment of exercise.
For instance, P13 expressed a greater appreciation for everyday physical activities: “My attitude going into this was, well, I don’t do a lot, I’m not doing enough. […] And this helped me understand that I actually am doing things. When I work out in the garden and I’m digging holes, […] I didn’t think that it was exercise or fitness in any way.” This increase in awareness led P13 to experience a gradual shift in confidence: “Each day I got a little bit more confident. […] It wasn’t a light bulb moment. It was a small, everyday, slow build.” P27 described adopting a more self-compassionate stance toward “bad days,” attributing this shift to Bloom’s flexibility: “Maybe this day wasn’t a good day. It’s fine because I have another day. […] if it’s not a good day, and I want to have a light workout, [Bloom] gives me that option. So I think that really helped me get motivated.” Similarly, P34 linked increased awareness of movement directly to their interactions with Beebo: “It definitely has changed the way I think about it […] if I go into the kitchen or outside or something, I’m doing all these additional steps. […] I’m moving a whole lot more than I ever did. And I think it’s because of my little Beebo.” Lastly, some participants described an increased intrinsic enjoyment of PA, such as P1: “At first I was like, working out sucks. But now I’m kind of like, oh, no. It’s kind of just part of what I do and I enjoy. I look forward to working out now.”
@@ 6.2. Survey Data
Figure 7. Mindset-related survey outcomes. We plot mindset-related survey items for pre/post (top), weekly (middle), and daily (bottom) survey responses. Scores represent group means on a 5-point Likert scale, with higher values reflecting more positive orientations toward PA. For daily and weekly surveys, differences ($\Delta$) are derived from linear regression coefficients and therefore may not equal the arithmetic difference between the displayed pre- and post-study means. Across various mindset-related survey measures, treatment participants showed greater increases than control participants.
Mindset-related survey outcomes. This figure plots survey outcomes from pre/post (top), weekly (middle), and daily (bottom) surveys as horizontal arrows. Green arrows represent the treatment group and gray arrows the control group. Across measures, including satisfaction with physical health and activity level, perceptions of exercise, and self-reported activity identity, treatment participants showed greater increases over time compared to controls.
Next, we report on survey outcomes organized by topic: physical activity, mindset and satisfaction, usability, and user experience. As discussed in Section 5.5.2, no statistical significance testing was conducted and findings should be interpreted as descriptive patterns of change. Full survey results are provided in Supplementary Tables 11 (pre/post), 12 (daily), and 13 (weekly).
@@ 6.2.1. Physical Activity
Self-reported levels of PA, assessed via the International Physical Activity Questionnaire (IPAQ), increased substantially in both groups with similar magnitude (+334+334 MET-min in the treatment group vs. +345+345 MET-min in the control group). Exercise Stage of Change (1 = pre-contemplation, 5 = action) advanced on average by +1.0+1.0 stages in the treatment group and by +0.7+0.7 in the control group.
@@ 6.2.2. Mindset & Satisfaction
Mindset-related outcomes exhibited the most prominent treatment-control differences, as shown in Figure 7. Treatment participants reported a larger increase in PA adequacy mindset, namely the belief that PA is beneficial to their health (+1.2+1.2 points from pre- to post-study vs. +0.8+0.8 in control; 5-point Likert scale), while both groups showed a similar increase in process mindset, the degree to which exercise is seen as inherently enjoyable (+0.3+0.3 vs. +0.2+0.2 points). Satisfaction with PA levels showed larger increases in the treatment group (+1.4+1.4 vs. +1.0+1.0 points), as did satisfaction with physical health (+0.9+0.9 vs. +0.6+0.6 points). Meanwhile, perceived barriers to activity declined slightly in both groups (0.2-0.2 vs. 0.3-0.3 points), while self-efficacy was stable ($0.0$ vs. $0.0$ points), suggesting that the groups differed in their beliefs about the benefits and enjoyment of PA, but not in their perceived barriers or efficacy.
Weekly and daily surveys exhibited similar trends. We report daily and weekly linear regression slopes multiplied by the 28-day or four-week study duration (respectively), producing an estimate of total change over the study period, all on a 5-point Likert scale. Daily survey results indicate that treatment participants showed greater increases in satisfaction with their current PA levels (+0.6+0.6 vs. +0.2+0.2 points), identifying as a physically active person (+0.8+0.8 vs. +0.3+0.3 points), and hopefulness about their health (+0.1+0.1 vs. 0.1-0.1 points). Weekly survey results similarly indicate that treatment participants showed greater increases in perceptions that exercising is easy (+0.6+0.6 vs. +0.1+0.1 points; on a scale from 1: difficult to 5: easy).
6.2.3. Usability
@@
H2 Treatment-control difference in daily step count change from baseline to study period -692 (432)
H3 Treatment-control difference in the weekly rate of change ($\Delta$ steps/day per study week) +229 (137)
@@
Burned (kcal) H2 Treatment-control difference in daily kcal burned change from baseline to study period -7.09 (31.3)
H3 Treatment-control difference in the weekly rate of change ($\Delta$ kcal/day per study week) +7.80 (9.82)
@@
H2 Treatment-control difference in daily exercise min change from baseline to study period -2.08 (4.95)
H3 Treatment-control difference in the weekly rate of change ($\Delta$ min/day per study week) +1.51 (1.57)
@@
Running (km) H2 Treatment-control difference in daily distance walked/run (km) change from baseline to study period -0.239 (0.198)
H3 Treatment-control difference in the weekly rate of change ($\Delta$ km/day per study week) +0.071 (0.063)
Table 7. Wearable Data (Quantitative) Results. We report mean (SE) parameter estimates from our model for each hypothesis. *** denotes $p<0.001$ and ** denotes $p<0.01$, where $p$-values are Holm-adjusted for multiple comparisons. Significant results are bolded.
@@ 6.2.4. User Experience
Participants in the treatment group consistently rated the quality of advice and overall user experience more favorably than the control group at both pre- and post-study measurements. We replicate Jörke et al.’s (Jörke et al., 2025) strong initial impressions of the GPTCoach system (pre-study mean of $4.4$ in treatment vs. $4.1$ in control; 5-point Likert scale). As with usability, ratings declined pre- to post-study in both conditions (0.4-0.4 vs. 0.5-0.5 points).
Treatment participants rated shared decision-making more highly at both pre- and post-study measurements, with scores increasing pre/post in treatment while decreasing in control (+0.1+0.1 vs. 0.2-0.2; 5-point Likert scale), suggesting that goal setting was perceived as a more collaborative activity in the treatment group. The treatment group also reported greater increases in insights into their personal data, measured via the Technology Support Reflection Inventory (TSRI) than the control group (+0.1+0.1 vs. 0.3-0.3; 5-point Likert scale). In addition, treatment participants also rated their interactions as more human-like (post-study mean of $3.5$ vs. $2.8$).
6.3. Wearable Data
We now turn to objective measures of PA outcomes, complementing the self-reported survey outcomes. Table 7 presents coefficient means, standard errors, and significance levels for each hypothesis from our mixed-effects model. Across all four outcomes—step count, active energy burned (kcal), exercise time (min), and distance walking/running (km)—mean PA during the four-week study period was significantly higher than mean baseline PA (H1), corresponding to moderate-to-large standardized effects (Cohen’s d=0.59d=0.59-$0.75$). On average, participants walked 1,680 more steps/day ($p<0.001$, $d=0.75$), burned 87.1 more kcal/day ($p<0.01$, $d=0.59$), spent 13.2 more minutes/day exercising ($p<0.01$, $d=0.60$), and walked 0.76 km/day more ($p<0.001$, $d=0.74$) during the study period than at baseline. These effects translate to substantial increases in activity, doubling the proportion of participants meeting the recommended 150 min/week of PA (for Disease Control and Prevention, 2022), increasing from 36% at baseline to 72% during the study (41% $\rightarrow$ 71% in control; 31% $\rightarrow$ 74% in treatment).
Our between-condition hypotheses were not statistically significant (H2: p=0.78p=0.78-$1.0$, H3: p=0.77p=0.77-$1.0$), and given the small observed effect sizes (d=0.22d=-0.220.230.23) and the study’s limited power to detect effects of this magnitude, the results are statistically inconclusive. However, across all four PA outcomes, we observed consistent directional patterns. For H2 (difference in mean PA), coefficients were negative for every outcome, indicating that the control group exhibited higher average PA levels than the treatment group during the study period. For H3 (differences in weekly rate of change), coefficients were positive for every outcome, indicating that activity in the treatment group declined less or increased slightly over the four weeks relative to control.
In summary, H1 was confirmed: both groups showed large and significant increases in PA over the baseline. H2 and H3 were not statistically conclusive, but the consistent directional patterns suggest that the control group exhibited a larger initial increase in PA levels followed by a steeper week-to-week decline, whereas the treatment group showed a smaller initial increase but more stable PA persistence. However, we caution that these patterns should be viewed as hypothesis generating rather than confirmatory.